A library has 8 tables and 42 chairs. There are several small tables with 4 chairs each and some large tables with 6 chairs each. How many tables of each type are in the library?
Answer by stanbon(75887)

A library has 8 tables and 42 chairs. There are several small tables with 4 chairs each and some large tables with 6 chairs each. How many tables of each type are in the library?
---------
Table equation: S + L = 8
Chair equation:4S +6L = 42
-------------------------------
Modify for Elimination:
2S + 2L = 16
2S + 3L = 21
-------------------
Subtract and solve for "L":
L = 5 (# of large tables)
---
Solve for "S":
S + L = 8
S = 3 (# of small tables)
=================
Cheers,
Stan H.
================